For normal incidence of light from a medium of refractive index 1.42 onto medium of refractive index 1.64, calculate the reflection and transmission coefficients.
1
Expert's answer
2020-04-05T14:32:35-0400

The reflection coefficient


"R=\frac{(n_2-n_1)^2}{(n_2+n_1)^2}=\frac{(1.64-1.42)^2}{(1.64+1.42)^2}=0.0052"


The transmission coefficient


"T=1-R=\frac{4n_1n_2}{(n_2+n_1)^2}=\frac{4\cdot1.42\cdot1.64}{(1.64+1.42)^2}=0.9948"